Write a Python program/s to calculate you will need to create variables for each equation.

  1. 1+3
  2. 10 - 4
  3. 7 time 6
  4. 81 divided by 9
  5. 80 divided by 11
  6. 12 raised to the 5th power
  7. 10 times 2 + 80 -63 divided 16
  8. 365 time 6 divided by 52
  9. 32000.00 divded by 16.50
  10. 200.00 times 40 times 26
Answer #1

a=1+3
b=10-4
c=7*6
d=81/9
e=80/11
f=12**5
g=10*2+80-63/16
h=365*6/52
i=32000.00/16.50
j=200*40*26


// If any doubt please comment